40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 12 11:19:04 2010, kiwamu, Update, CDS, current status 
    Reply  Thu Aug 12 14:01:53 2010, josephb, Update, CDS, current status 
       Reply  Thu Aug 12 18:30:08 2010, kiwamu, Update, CDS, Re: current status 
    Reply  Tue Aug 17 02:07:11 2010, Yoichi, Update, CDS, DACs have correct number of channels 
Message ID: 3405     Entry time: Thu Aug 12 11:19:04 2010     Reply to this: 3408   3428
Author: kiwamu 
Type: Update 
Category: CDS 
Subject: current status 

Current status of the new CDS test (summaries can be found on the wiki)

- Cables 

  All the necessary cables are in our hands, such as SCSIs, 37 pin D-sub and 3 pin power cables.

  With a big help from Jenne and Koji, we made 37 pin F/M cables and 3 pin power cables on the last Tuesday.

  Now all the cables are connected to the machines except for the 3 pin power cables.

  To install the power cables I will switch off Sorensens at new 1X5  for a minute.

 

- Suspension model file 

  The model file named C1SUS was made and I succeeded in compiling and building it. So it is ready for the test.

  But some medm screens look like not correctly complied.

  For example the channel names which are listed on C1SUS_MONITOR_ADC0.adl aren't correctly assigned.

 

- ADCs 

 All the ADC channels are working. Also the channel assignments are correct.

 I checked all the ADC channels if it was working while I ran the front end realtime code C1SUS.

 By injecting a signal from a function generator directly to the SCSIs, I could clearly see the numbers hopping on medm screens.

 

- DACs 

  None of them are working although the computer can recognize that all three DAC cards are mounted.

  I think something in the IOP file and the model file may be wrong because this symptom looks similar to that of C1ISCEX we tested two weeks before ( see this entry ) 

  I have to fix it anyhow because the DACs are very necessary part for this damping test.

 

- Binary Outputs

  They didn't work at all. Even the computer didn't recognize the binary output cards.

  I think I should put some magical components on the model files.

 

- Conversion of the filter coefficients 

  The conversion of the filter coefficients has been doen yesterday.

  It looks running well because I can load and unload these filters on medm screens.

  I manually copied the filter coefficients from the current suspension filter file to that of new filter file.

 The current file can be found under /cvs/cds/caltech/chans/,  and the new one is under /cvs/cds/rtcds/caltech/c1/chans


Suspended works

   - Binary Outputs

   - simlink realtime model with the new CDS PARTS ( see the notice from Joe )

 

To be done

   - let DACs work

   - damping tests

 

ELOG V3.1.3-